how DO YOU add, subtract, multiply, divide and simplify fractions, as well as how to convert decimal to numeric fractions.?
Answer by TutorDelphia(193)

Add fractions: make sure the denominator is the same, add the tops
---
Subtract fractions: make sure the denominator is the same, subtract the tops
---
Multiply: top time top, bottom times bottom
---
Divide: Dividing fractions is as easy as pie, all you do is flip the second and
multiply (take the reciprocal of the number you are dividing by and multiply)
---
Simplify fractions: find the largest number that goes into both the denominator and numerator and divide both numbers by that number: ex 4/16 ...4 is the biggest that goes into both 1/4
---
Convert decimals to fractions:
.1 = 1/10
.12 = 12/100
.01 = 1/100
.001= 1/1000
.123=123/1000